# The Reliability of Remotely Piloted Aircraft System Performance under Communication Loss and Latency Uncertainties

### v1

@@ 3.1 RCP Concept
In RCP, a transaction is the process of the human uses to give clearance, order commands, and send flight information, and is completed based on subjective confidence. Moreover, the transaction time include communication steps such as the speaking time (e.g., averaged 2.6 seconds for pilot and 3.8 seconds for controllers (Sollenberger et al., 2003)) and the reaction time of the controller and the pilot (i.e., 1 second), as well as the Air-to-Ground and Ground-to-Ground signal transmission time. Notably, the European Organisation for Civil Aviation Equipment (EUROCAE) ED136 provides the acceptable maximum latency of $100ms$ from the air traffic controller to the ground stations, $2ms$ for A/G communication, and $50ms$ for radio activation (ICAO ACP-WGF24 WP15, 2011). In the case of multiple transactions, the TT is the time to complete the most stringent transaction. RCP type is the specific label defines the communication transaction performance standard, denoting values for transaction time, availability, continuity and integrity. Five RCP types are prescribed (RCP10, RCP60, RCP120, RCP240, RCP400) (ICAO, 2006). RCP10 standard requires the communication transaction is likely to be finished within 10 seconds, where 10 seconds is the transaction time. It is mostly used for enhancing the aeronautical horizontal separation violation (i.e., 5nm) intervention capability of controller. Continuity can be critical for RCP10, where the transaction is likely to be finished within 10 seconds without the time to repeat the voice messages. To meet RCP10, the transaction should be completed with a single attempt. RCP60 usually complements RCP10 with data communication (i.e., data link). RCP120 and RCP240 are used for supporting separation violation intervention capability in the extended range of 15nm radius or 30/30 (30nm lateral and 30nm longitudinal). RCP240 is the mandatory for oceanic airspace of application with certain separation standards as noted by the FAA (The Performance Based Operations Aviation Rulemaking Committee (2018), PARC). Beyond 30/30 separation, RCP400 is the standard for SATVOICE and SATCOM where communication services are provided through HF signal. A recent collaborative study between the FAA and JetBlue Airways conducted in Fall 2017 validated the onboard SATVOICE system is able to satisfy the RCP400 criteria and SATVOICE is viable for RCP400 (The Performance Based Operations Aviation Rulemaking Committee (2018), PARC). The communication equipment carried by the aircraft should at least satisfy the appropriate required RCP type to receive approval of operations. Other RCP types may also be established based on experience and usage. The development of RCP concept enables regulatory authorities to grant operational approvals for specific areas based on the capability to meet a specified RCP, rather than solely on the installation or carriage of specific communications equipment. The airspace authority prescribes the RCP for a given area of airspace or a specific procedure after determining RCP types, as well as maintaining compliance of RCP among different airspace users.
@@ 3.2 Mathematical Formulations
We start by defining $\tau_{msg}$ as the duration of the message to be transmitted in the communication system (i.e., Transaction Time). We further assume that the signal transition between on and off states follows the Continuous-Time Markov Chain (CTMC), with exponential waiting times for each state. That is,
@@ 3.2.1 Availability
[…]
@@ 3.2.2 Continuity
$P_{C}(\tau_{msg})$ $=P(X(\tau_{0})=1,X(\tau_{1})=1,\ldots,X(\tau_{n})=1)$ (16)
$=P(X(\tau_{0})=1)\cdot P(X(\tau_{1})=1|X(\tau_{0})=1)\cdot P(X(\tau_{2})=1|\{X(\tau_{0})=1\cap X(\tau_{1})=1\}),$ (17)
$\ldots,P(X(\tau_{n})=1|\{X(\tau_{0})=1\cap X(\tau_{1})=1,\ldots,\cap X(\tau_{n-1})=1\})$ (18)
@@
$P_{C}(\tau_{msg})$ $=P(X(\tau_{0})=1)\cdot P(X(\tau_{1})=1|X(\tau_{0})=1),\ldots,P(X(\tau_{n})=1|X(\tau_{n-1})=1))$ (19)
$=P(X(\tau_{0})=1)\cdot\Pi_{k=1}^{n}P(X(\tau_{k})=1|X(\tau_{k-1})=1))$ (20)
$=P_{A}(\tau_{0})\cdot\Pi_{k=1}^{n}P(X(\tau_{k})=1|X(\tau_{k-1})=1))$ (21)
@@ 3.3 Communicability
[…]
@@ 4 Flight Simulator
In this section, we briefly introduce the 6DOF flight simulation environment used in this research. We adopt the autonomous aerospace verification challenge benchmark, which was released as the basis for analyzing the detailed behavior of autonomous maneuvers for a remotely piloted high performance F-16 Falcon (Heidlauf et al., 2018).
The benchmark model is based on the textbook model, where the performance of F-16 has been extensively studied since the 1970s (Nguyen, 1979; Seto et al., 1999; Stevens et al., 2015). The benchmark uses the linearized version of the nonlinear flight dynamics of the aircraft about a steady-state trim point (i.e., without gain scheduling), by assuming the autonomous aircraft is operating without sideslip, on a flat earth using the north-east-down frame, and with approximated aerodynamic coefficients. The benchmark model further ignores the leading edge flaps and experimental thrust vectoring control (Sonneveldt, 2006). The aerodynamic coefficients are reasonable approximations for the subsonic flight regime. Euler angle based equations of motion are used, however, they are limited to moderate intensity maneuvers, as they will encounter singularities when the pitch angle $\theta$ exceed ±$\pm\frac{\pi}{2}$, in which case quaternions should be used (Sonneveldt, 2006). Further explanation of the equations of motion is given in Section 4.1.
@@ 4.1 Flight Dynamics of the Simulator
The relationship between the body frame velocities $[u,v,w]^{T}$ and the Earth frame velocities $[\dot{x}_{e},\dot{y}_{e},\dot{h}_{e}]^{T}$ can be written as the product of three rotation matrices:
@@
$=\begin{bmatrix}\cos\psi&-\sin\psi&0\\ \sin\psi&\cos\psi&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}\cos\theta&0&\sin\theta\\ 0&1&0\\ -\sin\theta&0&\cos\theta\end{bmatrix}\begin{bmatrix}1&0&0\\ 0&\cos\phi&-\sin\phi\\ 0&\sin\phi&\cos\phi\end{bmatrix}\begin{bmatrix}u\\ v\\ w\end{bmatrix}$ (45)
$=\begin{bmatrix}\cos\theta\cos\psi&\cos\theta\sin\psi&-\sin\theta\\ \sin\phi\sin\theta\cos\psi-\cos\phi\sin\psi&\sin\phi\sin\theta\sin\psi+\cos\phi\cos\psi&\sin\phi\cos\theta\\ \cos\phi\sin\theta\cos\psi+\sin\phi\sin\psi&\cos\phi\sin\theta\sin\psi-\sin\phi\cos\psi&\cos\phi\cos\theta\end{bmatrix}\begin{bmatrix}u\\ v\\ w\end{bmatrix}$ (46)
@@ 4.2 Controller Design
[…]
is desired, where $Q$ and $R$ are penalty metrics in symmetric positive-semidefinite forms (i.e., $\mathcal{J}\geq 0$). The optimal solution can be solved with Riccati equation, or gradient-based approach when an analytical solution is infeasible (Choi and Seo, 1999; Ashraf et al., 2018; Priyambodo et al., 2020).
@@ 4.2.2 Lower-Level Control
[…]
@@
Figure 5: Aircraft trajectory in the vertical plane when comparing $P_{A}=1$ and $P_{A}=0.95$ with no latency in both cases. For a given $P_{A}=0.95$, we first generate the signal state masks square wave for the entire simulation duration Figure 2. Then, apply to the mask to $\mathbf{u}(t)$ to inference the state derivatives.
Following the formulation in Section 3.2, we simulate communication C2 link loss with CTMC between the on and off states Figure 6. The overhead aircraft trajectory from two simulation runs with $P_{A}=1$ and $P_{A}=0.95$ is also shown Figure 5, along with a comparison of higher-level control inputs (Figure 8) and lower-level control inputs (Figure 9). The mission is defined as a waypoint following task, where the RPAS is commanded to reach a minimum slant range to each waypoint before proceeding to the next. The RPAS starts from the origin, while the airspace of interest is $[-12,000,4,000]$ feet on the east/west direction, and $[0,20,000]$ feet to the north/south direction. The altitude range is around $[3,800,4,600]$ feet with Vt between $[535,555]$ feet/second (Figure 7). We choose to regulate first three higher-level reference values $\mathbf{r}=[N_{z}^{\text{cmd}},p^{\text{cmd}},N_{yr}^{\text{cmd}},\delta_{t}^{\text{cmd}}]$ to 00 once the C2 link is lost. The objective of regulating these reference values to zero is to make sure the F-16 can return to a steady straight and level flight once the RP connection is lost.
@@ 5.1 Definition of Mission Success
The mission is defined as the waypoint following task in the 3D airspace. Following the notations and derivations used for HLC in Section 4.2.1, we further compute the slant range $R$ by the $L^{2}$ distance between the current waypoint and the current position (e.g., $R=\sqrt{\Delta_{x}^{2}+\Delta_{y}^{2}+\Delta_{z}^{2}}$). The simulation checks the slant range to each current waypoint, and switches to the next waypoint only if the condition $R<R_{\text{threshold}}$ is met. Once all waypoints are reached, the simulation changes the mode to Done, and sets $\psi^{\text{cmd}}$ to 00 in HLC to maintain a level flight. Mission completion is defined as the RPAS reaches all given waypoint in the current scenario with a Done status from the simulator. Specifically, we select two waypoint scenarios to represent a hard case and a simple case as in Figure 11. Figure 11(a) is considered the hard case with 4 waypoints and requires a minimum radius turn of RPAS, while Figure 11(b) is the simple scenario.
@@ 5.2 Mission Success Envelope
The simulation was completed on a high-performance workstation wit Intel Xeon w9-3495X processor with 56 cores, 112 threads, and a maximum turbo frequency of up to 4.8 GHz. We collect slightly over 1 million Monte Carlo Simulation samples for each scenario, utilizing the multiprocessing tools (Aziz et al., 2021). We further store the Boolean indicating whether the waypoint following mission is completed, and the mission completion time if completed.
@@ 5.2.1 Mission Success Rate
Mission success rate is a straightforward performance indicator. It quantifies how frequently an operation meets the intended goals, thereby informing engineers or decision-makers on overall system reliability expectations, and the regions in need of improvement (Chamberlain, 1966; Jacklin, 2019; Zhao et al., 2020). In this work, mission success rate is calculated as the ratio between completed simulations and total number of simulations for any given parameter pair $\varepsilon$ and $P_{A}$. It stands for the percentage or proportion of attempts that meet the predefine success criteria given above.
@@ 5.3 Analysis of Communicability
As introduced in Section 3.3, communicability as a combined metric to represent RCP key metrics, considers both $\varepsilon$ and $P_{A}$ for a given message transaction time $\tau_{msg}$. $\tau_{msg}$ is computed based on the message size (in bits) and the transmission bitrate (in bits per second) as,
@@
where $\tau_{msg}$ is the message transmission time (seconds), $S_{bits}$ is the size of the message in bits, and $R_{bitrate}$ is the transmission bitrate in bits per second (bps). In this simulation, seven double float control inputs are continuous transmitted between the RPA and RPS, corresponding to a total of $7\times 8=56$ bytes, or,
@@
$P_{\text{comm}}$ is then calculated with Equation 36. We further draw the mission success rate and mission completion time versus $P_{\text{comm}}$ for $\varepsilon$ intervals for both scenarios, as in Figure 17. Communicability analysis facilitates better understanding of the maximum tolerable latency value for a given data transmission system. For instance, it is obvious that in Figure 17(b), latency of $0.08$s still leads to high mission success rate for high communicability in simpler scenarios. Figure 17(a) also shows the impact of latency blockage where a higher latency (i.e., $\varepsilon\in(0.07,0.08]$) can achieve higher mission success rate than lower latency (i.e., $\varepsilon\in(0.06,0.07]$), even with high communicability (i.e., $P_{\text{Comm}}=1$). Other aviation communication system (i.e.,Ka/Ku-band/L-band SATCOM, Aeronautical Mobile Airport Communication System (AeroMACS)) with different bitrate (i.e., AeroMACS maxed 9.2 Mbps.) can be evaluated in a similar way.

### v2

@@ 3.1 The RCP Concept
In RCP, a transaction is the process of the human uses to give clearance, order commands, and send flight information, and is completed based on subjective confidence. Moreover, the transaction time include communication steps such as the speaking time (e.g., averaged 2.6 seconds for pilot and 3.8 seconds for controllers (Sollenberger et al., 2003)) and the reaction time of the controller and the pilot (i.e., 1 second), as well as the Air-to-Ground and Ground-to-Ground signal transmission time. Notably, the European Organisation for Civil Aviation Equipment (EUROCAE) ED136 provides the acceptable maximum latency of $100ms$ from the air traffic controller to the ground stations, $2ms$ for A/G communication, and $50ms$ for radio activation (ICAO ACP-WGF24 WP15, 2011). In the case of multiple transactions, the TT is the time to complete the most stringent transaction. RCP type is the specific label defines the communication transaction performance standard, denoting values for transaction time, availability, continuity and integrity. Five RCP types are prescribed (RCP10, RCP60, RCP120, RCP240, RCP400) (ICAO, 2006). RCP10 standard requires the communication transaction is likely to be finished within 10 seconds, where 10 seconds is the transaction time. It is mostly used for enhancing the aeronautical horizontal separation violation (i.e., 5nm) intervention capability of controller. Continuity can be critical for RCP10, where the transaction is likely to be finished within 10 seconds without the time to repeat the voice messages. To meet RCP10, the transaction should be completed with a single attempt. RCP60 usually complements RCP10 with data communication (i.e., data link). RCP120 and RCP240 are used for supporting separation violation intervention capability in the extended range of 15nm radius or 30/30 (30nm lateral and 30nm longitudinal). RCP240 is the mandatory for oceanic airspace of application with certain separation standards as noted by the FAA (The Performance Based Operations Aviation Rulemaking Committee (PARC), 2018). Beyond 30/30 separation, RCP400 is the standard for SATVOICE and SATCOM where communication services are provided through HF signal. A recent collaborative study between the FAA and JetBlue Airways conducted in Fall 2017 validated the onboard SATVOICE system is able to satisfy the RCP400 criteria and SATVOICE is viable for RCP400 (The Performance Based Operations Aviation Rulemaking Committee (PARC), 2018). The communication equipment carried by the aircraft should at least satisfy the appropriate required RCP type to receive approval of operations. Other RCP types may also be established based on experience and usage. The development of RCP concept enables regulatory authorities to grant operational approvals for specific areas based on the capability to meet a specified RCP, rather than solely on the installation or carriage of specific communications equipment. The airspace authority prescribes the RCP for a given area of airspace or a specific procedure after determining RCP types, as well as maintaining compliance of RCP among different airspace users.
@@ 3.2 Mathematical Formulations
We start by defining $\tau_{msg}$ as the duration of the message to be transmitted in the communication system (i.e., Transaction Time). We further assume that the signal transition between on and off states follows the Continuous-Time Markov Chain (CTMC), with exponential waiting times for each state. Markov-based methods have been widely adopted in reliability engineering for modeling dynamic systems. For example, Markov models have been combined with event-tree/fault-tree analysis for dynamic reliability assessment (Bucci et al., 2008), Petri nets with aging tokens have been used for modeling degrading and repairable systems (Volovoi, 2004), discrete-time Bayesian networks have been applied to systems with common cause failures (Guo et al., 2021b), and Dynamic Object Oriented Bayesian Networks have been formalized for complex system reliability evaluation (Weber and Jouffe, 2006). That is,
@@ 3.2.1 Availability
$\frac{dP_{A}(t)}{dt}$ $=\text{inflow rate}-\text{outflow rate}$ (4)
@@ 3.2.2 Continuity
$P_{C}(\tau_{msg})$ $=P\bigl(X(\tau_{0})=1,X(\tau_{1})=1,\ldots,X(\tau_{n})=1\bigr)$
$=P\bigl(X(\tau_{0})=1\bigr)\cdot P\bigl(X(\tau_{1})=1\mid X(\tau_{0})=1\bigr)$
$\quad\cdot P\bigl(X(\tau_{2})=1\mid X(\tau_{0})=1\cap X(\tau_{1})=1\bigr)\cdots$
$\quad\cdot P\bigl(X(\tau_{n})=1\mid$ $X(\tau_{0})=1\cap X(\tau_{1})=1\cap\cdots\cap X(\tau_{n-1})=1\bigr)$ (16)
@@
$P_{C}(\tau_{msg})$ $=P\bigl(X(\tau_{0})=1\bigr)\cdot P\bigl(X(\tau_{1})=1\mid X(\tau_{0})=1\bigr)\cdots$
$\quad\cdot P\bigl(X(\tau_{n})=1\mid X(\tau_{n-1})=1\bigr)$
$=P\bigl(X(\tau_{0})=1\bigr)\cdot\prod_{k=1}^{n}P\bigl(X(\tau_{k})=1\mid X(\tau_{k-1})=1\bigr)$
$=P_{A}(\tau_{0})\cdot\prod_{k=1}^{n}P\bigl(X(\tau_{k})=1\mid X(\tau_{k-1})=1\bigr)$ (17)
@@ 3.3 Communicability
$P_{\text{comm}}$ $=P_{A}\cdot\bigg[\frac{\int_{\tau_{msg}+\varepsilon}^{\infty}(\tau-(\tau_{msg}+\varepsilon))\cdot f(\tau)d\tau}{\int_{0}^{\infty}\tau\cdot f(\tau)d\tau}\bigg]$ (30)
@@ 4 Flight Simulator
In this section, we briefly introduce the 6DOF flight simulation environment used in this research. Flight dynamics simulation has long served as a key tool for reliability and safety verification of aerospace systems, from early applications such as the Dynamic Flowgraph Methodology demonstrated on the Titan II space launch vehicle digital flight control system (Yau et al., 1995), to recent risk-based control system designs for autonomous vehicles verified through formal methods and simulation-based testing (Johansen et al., 2023). We adopt the autonomous aerospace verification challenge benchmark, which was released as the basis for analyzing the detailed behavior of autonomous maneuvers for a remotely piloted high performance F-16 Falcon (Heidlauf et al., 2018).
The benchmark model is based on the textbook model, where the performance of F-16 has been extensively studied since the 1970s (Nguyen, 1979; Seto et al., 1999; Stevens et al., 2015). The benchmark uses the linearized version of the nonlinear flight dynamics of the aircraft about a steady-state trim point (i.e., without gain scheduling), by assuming the autonomous aircraft is operating without sideslip, on a flat earth using the north-east-down frame, and with approximated aerodynamic coefficients. The benchmark model further ignores the leading edge flaps and experimental thrust vectoring control (Sonneveldt, 2006). The aerodynamic coefficients are reasonable approximations for the subsonic flight regime. Euler angle based equations of motion are used, however, they are limited to moderate intensity maneuvers, as they will encounter singularities when the pitch angle $\theta$ exceed $\pm\frac{\pi}{2}$, in which case quaternions should be used (Sonneveldt, 2006). Further explanation of the equations of motion is given in Section 4.1.
@@ 4.1 Flight Dynamics of the Simulator
The relationship between the body frame velocities $u,v,w]^{T}$ and the Earth frame velocities $\dot{x}_{e},\dot{y}_{e},\dot{h}_{e}]^{T}$ can be written as the product of three rotation matrices:
@@
where the product of the three rotation matrices yields the Direction Cosine Matrix (DCM):
$R_{y}R_{p}R_{r}=\begin{bmatrix}\cos\psi&-\sin\psi&0\\ \sin\psi&\cos\psi&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}\cos\theta&0&\sin\theta\\ 0&1&0\\ -\sin\theta&0&\cos\theta\end{bmatrix}\begin{bmatrix}1&0&0\\ 0&\cos\phi&-\sin\phi\\ 0&\sin\phi&\cos\phi\end{bmatrix}$
$\begin{bmatrix}\cos\theta\cos\psi&\cos\theta\sin\psi&-\sin\theta\\ \sin\phi\sin\theta\cos\psi-\cos\phi\sin\psi&\sin\phi\sin\theta\sin\psi+\cos\phi\cos\psi&\sin\phi\cos\theta\\ \cos\phi\sin\theta\cos\psi+\sin\phi\sin\psi&\cos\phi\sin\theta\sin\psi-\sin\phi\cos\psi&\cos\phi\cos\theta\end{bmatrix}$ (42)
so that $\dot{x}_{e},\dot{y}_{e},\dot{z}_{e}]^{T}=R_{y}R_{p}R_{r}\,[u,v,w]^{T}$.
@@ 4.2 Controller Design
$\min\mathcal{J}=\frac{1}{2}\int_{0}^{\infty}\left(\mathbf{x}(t)^{T}Q\mathbf{x}(t)+\mathbf{u}(t)^{T}R\mathbf{u}(t)\right)dt$ (60)
is desired, where $Q$ and $R$ are penalty metrics in symmetric positive-semidefinite forms (i.e., $\mathcal{J}\geq 0$). The optimal solution can be solved with Riccati equation, or gradient-based approach when an analytical solution is infeasible (Choi and Seo, 1999; Ashraf et al., 2018; Priyambodo et al., 2020). In the context of reliability-constrained control, LQR has also been extended to incorporate interval uncertainty and nonprobabilistic time-dependent reliability constraints for satellite attitude-vibration control (Yang et al., 2023).
@@ 4.2.2 Lower-Level Control
For longitudinal motion, the short-period approximation to the aircraft dynamics is linearized around the nominal flight condition, as referenced in Table 3.6-3 of (Stevens et al., 2015). The states of interests are $\mathbf{x}_{\text{long}}=[\alpha,q,\int e_{N_{z}}dt]^{T}$, and $\mathbf{y}_{\text{long}}=[\alpha,q,N_{z}]^{T}$, where $\int e_{N_{z}}dt$ is the integral of the normal acceleration error. $e_{N_{z}}=N^{ref}_{z}-N_{z}$ is the difference between the reference normal acceleration from the Autopilot and the measured normal acceleration. The control input for longitudinal motion is the elevator deflection in radians $\mathbf{u}_{\text{long}}=[\delta_{e}]$.
For lateral motion, a separated LQR is introduced with the state variables $\mathbf{x}_{\text{lat}}=[\beta,p,r,\int e_{\psi}dt,\int e_{N_{yr}}dt]^{T}$, and $\mathbf{y}_{\text{lat}}=[\beta,p,r,\psi,N_{yr}]^{T}$ with control inputs $\mathbf{u}_{\text{lat}}=[\delta_{a},\delta_{r}]^{T}$. Similar to the longitudinal controller, the lateral controller uses state feedback combined with error integrals to ensure accurate tracking and disturbance rejection. The control inputs are computed as a linear combination of the current states and the integral of the error states.
@@
Figure 5: Aircraft trajectory in the vertical plane when comparing $P_{A}=1$ and $P_{A}=0.95$ with no latency in both cases. For a given $P_{A}=0.95$, we first generate the signal state masks square wave for the entire simulation duration as in Figure 6. Then, apply to the mask to $\mathbf{u}(t)$ to inference the state derivatives.
Following the formulation in Section 3.2, we simulate communication C2 link loss with CTMC between the on and off states Figure 6. The overhead aircraft trajectory from two simulation runs with $P_{A}=1$ and $P_{A}=0.95$ is also shown Figure 5, along with a comparison of higher-level control inputs (Figure 8) and lower-level control inputs (Figure 9). The mission is defined as a waypoint following task, where the RPAS is commanded to reach a minimum slant range to each waypoint before proceeding to the next. The RPAS starts from the origin, while the airspace of interest is [12,000,4,000][-12,000,4,000] feet on the east/west direction, and $[0,20,000]$ feet to the north/south direction. The altitude range is around $[3,800,4,600]$ feet with Vt between $[535,555]$ feet/second (Figure 7). We choose to regulate first three higher-level reference values $\mathbf{r}=[N_{z}^{\text{cmd}},p^{\text{cmd}},N_{yr}^{\text{cmd}},\delta_{t}^{\text{cmd}}]$ to 0 once the C2 link is lost. The objective of regulating these reference values to zero is to make sure the F-16 can return to a steady straight and level flight once the RP connection is lost.
@@ 5.1 Definition of Mission Success
The mission is defined as the waypoint following task in the 3D airspace. Following the notations and derivations used for HLC in Section 4.2.1, we further compute the slant range $R$ by the $L^{2}$ distance between the current waypoint and the current position (e.g., $R=\sqrt{\Delta_{x}^{2}+\Delta_{y}^{2}+\Delta_{z}^{2}}$). The simulation checks the slant range to each current waypoint, and switches to the next waypoint only if the condition $R<R_{\text{threshold}}$ is met. Once all waypoints are reached, the simulation changes the mode to Done, and sets $\psi^{\text{cmd}}$ to 0 in HLC to maintain a level flight. Mission completion is defined as the RPAS reaches all given waypoint in the current scenario with a Done status from the simulator. Specifically, we select two waypoint scenarios to represent a hard case and a simple case as in Figure 11. Figure 11(a) is considered the hard case with 4 waypoints and requires a minimum radius turn of RPAS, while Figure 11(b) is the simple scenario.
@@ 5.2 Mission Success Envelope
The simulation was completed on a high-performance workstation with Intel Xeon w9-3495X processor of 56 cores, 112 threads, and a maximum turbo frequency of up to 4.8 GHz. We collect slightly over 1 million Monte Carlo Simulation samples for each scenario, utilizing the multiprocessing tools (Aziz et al., 2021). We further store the Boolean indicating whether the waypoint following mission is completed, and the mission completion time if completed.
@@ 5.2.1 Mission Success Rate
Mission success rate is a straightforward performance indicator. It quantifies how frequently an operation meets the intended goals, thereby informing engineers or decision-makers on overall system reliability expectations, and the regions in need of improvement (Chamberlain, 1966; Jacklin, 2019; Zhao et al., 2020). Related formulations for phased-mission systems (PMS), commonly found in aerospace applications, have been studied using Semi-Markov Processes for non-exponential component lifetimes (Li et al., 2018a) and Markov regenerative processes for missions subjected to random shocks (Li et al., 2018b). Moreover, Bayesian networks have been employed to evaluate complex multi-state system reliability under common cause failures and epistemic uncertainties (Mi et al., 2018), and to assess the reliability of redundant control systems (Cai et al., 2012). In this work, mission success rate is calculated as the ratio between completed simulations and total number of simulations for any given parameter pair $\varepsilon$ and $P_{A}$. It stands for the percentage or proportion of attempts that meet the predefine success criteria given above.
@@ 5.3 Analysis of Communicability
As introduced in Section 3.3, communicability, as a combined metric to represent RCP key metrics, considers both $\varepsilon$ and $P_{A}$ for a given message transaction time $\tau_{msg}$. $\tau_{msg}$ is computed based on the message size (in bits) and the transmission bitrate (in bits per second) as,
@@
where $\tau_{msg}$ is the message transmission time (seconds), $S_{bits}$ is the size of the message in bits, and $R_{bitrate}$ is the transmission bitrate in bits per second (bps). In this simulation, seven double float control inputs are continuous transmitted between the RPA and RPS, corresponding to a total of 56 bytes, or,
@@
$P_{\text{comm}}$ is then calculated with Equation 32. We further draw the mission success rate and mission completion time versus $P_{\text{comm}}$ for $\varepsilon$ intervals for both scenarios, as in Figure 17. Communicability analysis facilitates better understanding of the maximum tolerable latency value for a given data transmission system. For instance, it is obvious that in Figure 17(b), latency of $0.08$s still leads to high mission success rate for high communicability in simpler scenarios. Figure 17(a) also shows the impact of latency blockage where a higher latency (i.e., $\varepsilon\in(0.07,0.08]$) can achieve higher mission success rate than lower latency (i.e., $\varepsilon\in(0.06,0.07]$), even with high communicability (i.e., $P_{\text{Comm}}=1$). Other aviation communication system (i.e.,Ka/Ku-band/L-band SATCOM, Aeronautical Mobile Airport Communication System (AeroMACS)) with different bitrate (i.e., AeroMACS maxed at 9.2 Mbps) can be evaluated in a similar way.